# The Effect of Tornadic Supercell Thunderstorms on the Atmospheric Muon Flux

### v3

@@ V Conclusion
Here we have presented simulations of the atmospheric muon flux in the presence of a supercell thunderstorm with an active tornado. We find that such a storm would introduce directional variations in the atmospheric muon flux of approximately 1%. If a storm can be observed over the course of an hour, a 1000 m2 muon detector would likely be capable of measuring variations in the local pressure field over approximately 1 kilometer length scales. If a prior model for the pressure field associated with the tornado itself can be developed, this technique could additionally be used to identify the presence and location of an active tornado.
The 1000 m2 muon detector considered in this paper is smaller than several existing cosmic ray telescopes. Detectors covering an even smaller area would likely be able to do similar science, and would have the advantage of portability, though will have to contend with proportionally larger statistical variations. In either case, measurements of the atmospheric muon flux in the vicinity of supercell thunderstorms could provide valuable information about the pressure field over a large volume that cannot be obtained using existing methods.

### v5

@@ V Conclusion
Here we have presented simulations of the atmospheric muon flux in the presence of a simulated supercell thunderstorm with an active tornado. We find that such a storm would introduce directional variations in the atmospheric muon flux of approximately 1%. If a storm can be observed over the course of an hour, a 1000 m2 muon detector would likely be capable of measuring variations in the local density field over approximately 1 kilometer length scales. If a prior model for the density field associated with the tornado itself can be developed, this technique could additionally be used to identify the presence and location of an active tornado.
The 1000 m2 muon detector considered in this paper is smaller than several existing cosmic ray telescopes. Detectors covering an even smaller area would likely be able to do similar science, and would have the advantage of portability, though will have to contend with proportionally larger statistical variations. Additionally, supercell thunderstorms present a high degree of storm-to-storm variability, and while in this case the density field variations appear to be associated with tornado formation, this may not be the case for all storms. In any case, measurements of the atmospheric muon flux in the vicinity of supercell thunderstorms could provide valuable information about the density field over a large volume that cannot be obtained using existing methods.